OpenStudy (agl202):

Quick answer check... Write the equation of the line that passes through (3, −2) and has a slope of 4 in point-slope form. Answers:

rishavraj (rishavraj):

whts ur answer??

OpenStudy (agl202):

I think it's x+2=4(y-3)

rishavraj (rishavraj):

nah see the slope of line passing from (x1 , y1) and x2 , y2) is (y2 y1)/(x2 - x1)

OpenStudy (agl202):

Ya

OpenStudy (agl202):

So would it be y+2=4(x-3)?

rishavraj (rishavraj):

yup :)))

OpenStudy (agl202):

Thx. :)

rishavraj (rishavraj):

yw :))